OpenStudy (anonymous):

Three consecutive terms in the expansion of (1+x)^n have coefficients 120,210, and 252. Find n. BINOMIAL THEOREM QUESTION @amistre64

OpenStudy (freckles):

Should be a matter of solving the following: \[\left(\begin{matrix}n \\ k\end{matrix}\right)=120 \\ \left(\begin{matrix}n \\ k+1\end{matrix}\right)=210 \\ \left(\begin{matrix}n \\ k+2\end{matrix}\right)=252\]

OpenStudy (anonymous):

I dont know how to do matrices, is there another way of doing this?

OpenStudy (freckles):

I didn't any matrices...

OpenStudy (freckles):

\[\left(\begin{matrix}n \\ k\end{matrix}\right)=\frac{n!}{k!(n-k)!}\]

OpenStudy (anonymous):

oh okay, got it let me try to solve it

OpenStudy (anonymous):

okay, im lost now lol. how do i cancel the k's?

OpenStudy (anonymous):

i know at some point n! reaches (n-k) factorial where then they both cancel out) but how do i get cancel out the other k!?

OpenStudy (freckles):

\[\frac{n!}{k!(n-k-2)! (n-k-1)(n-k)}=120 \\ \frac{n!}{k!(n-k-2)!(k+1)(n-k-1)}=210 \\ \frac{n!}{k!(n-k-2)!(k+1)(k+2)}=252\]

OpenStudy (anonymous):

I don't think i've learned this yet. My teacher might go through this tomorrow.

OpenStudy (anonymous):

yea, sorry didnt have to do this question, wasn't assigned

OpenStudy (amistre64):

(n 0) (n 1) (n 2) ... (n k) (n+1 0) (n 0) + (n 1) (n 1) + (n 2) (n 0) + (n 1) + (n 1) + (n 2) just athought

OpenStudy (freckles):

\[(n-k-1)(n-k)120=(k+1)(n-k-1)210=(k+1)(k+2)252 \\ \] \[(n-k)120=(k+1)210 \\ (n-k-1)210=(k+2)252 \]

OpenStudy (freckles):

\[(n-k)210-210=252k+504 \\ \frac{(k+1)(210)}{120}(210)-210=252k+504\]

OpenStudy (freckles):

so you can solve for k

OpenStudy (freckles):

and then you could go back and solve for n

OpenStudy (anonymous):

okay, awsome thanks

OpenStudy (amistre64):

how about we simply let n=10 ...

OpenStudy (anonymous):

yea that what i got from trial and error

OpenStudy (anonymous):

I got n=9 by doing Freckles way

OpenStudy (freckles):

k=3 \[(n-3)120=(3+1)210 \\n-3=\frac{(3+1)210}{120}=\frac{4(21)}{12}=\frac{21}{3}=7 \\ n-3=7 \]

OpenStudy (freckles):

What did you get for k @M0j0jojo ?

OpenStudy (anonymous):

i got 3 aswell maybe i probably messed on calculating n, let me check again

OpenStudy (anonymous):

yea found my mistake. got n=10

OpenStudy (freckles):

Also there is a lot of mistakes we can make in this problem since there so many steps.

OpenStudy (freckles):

I like @amistre64 way of saying let's just make n=10.

OpenStudy (amistre64):

lol, it just seemed easier to me that way

OpenStudy (anonymous):

lol, that what i did cuz from memory i know 10c5=252 so i was like oh maybe n=10

OpenStudy (amistre64):

if the 10 row was too large ... i was going to say 8 and just try and error it into submission

OpenStudy (amistre64):

\[n!=\int_{0}^{\infty} x^ne^{-x}~dx\] wonder if this is useful

OpenStudy (anonymous):

o.o

OpenStudy (freckles):

\[\left(\begin{matrix}n \\ k+1\end{matrix}\right)=\frac{n-k}{k+1}\left(\begin{matrix}n \\ k\end{matrix}\right) \\ \left(\begin{matrix}n \\ k+2\end{matrix}\right)=\frac{n-k-1}{k+2} \left(\begin{matrix}n \\ k+1\end{matrix}\right)\]

OpenStudy (freckles):

These probably would have gave us better but not much better equations to solve

OpenStudy (freckles):

I'm interesting in the integral way if you know how to approach that @amistre64

OpenStudy (amistre64):

rolling it around, not sure if its useful tho :) i cant recall the gamma function for nuthin off the top of my head

OpenStudy (amistre64):

120 C 1 = 120 120 C 2 not= 210 nC2 = 120 120 = x(x-1)/2 0 = n^2 -n-240; x=16 16C2 = 120 16C3 = 560 not it nC3 = 120 n(n-1)(n-2) = 720 n(n-1)(n-2) - 720 = 0 n = 10 10C3 = 120 10C4 = 210 10C5 .... would prolly have been my ultimate approach

OpenStudy (anonymous):

It would of been easier with just listing out pasca'ls triangle